FOURTH HUNGARIAN COLLOQUIUM ON LIMIT THEOREMS OF PROBABILITY AND STATISTICS List of participants

Antos András MTA SZTAKI, Budapest, Lágymányosi u. 11., H - 1111, Hungary antos@inf.bme.hu C. Azevedo Departamento de Matematica, Campus de Gualtar, Universidade de Minho, 4710 Braga, Portugal cecilia@math.uminho.pt Baran Sándor KLTE Matem. és Inform. Intézet, Debrecen, Pf. 12., H - 4010, Hungary barans@math.klte.hu J. Beirlant Dept. of Math., Catholic Univ. Leuven, Celestijnenlaan 200B, B-3001 Heverlee, Belgium jan.beirlant@wis.kuleuven.ac.be Berkes István MTA Rényi A. MKI, Budapest, Reáltanoda u. 13-15., P.O. Box 127, H - 1364, Hungary berkes@math-inst.hu Ph. Berthet Université de Rennes 1, Institut Mathematique, Campus de Beaulieu, 35042 Rennes Cedex, France berthet@maths.univ-rennes1.fr R. Bhattacharya Department of Mathematics, Indiana University, Bloomington, IN 47405-5701, USA bhattach@indiana.edu M. Bloznelis Dept. of Mathematics, Vilnius University, Naugarduko 24, Vilnius 2006 Lithuania blozn@mathematik.uni-bielefeld.de S. Borovkova ITS, Mekelweg 4, 2628 CD Delft, The Netherlands borovkova@math.rug.nl R. Bradley Department of Mathematics, Indiana University, Bloomington, IN 47405, USA bradleyr@indiana.edu M.D. Burke Dept. of Math. and Stat., Univ. of Calgary, 2500 University Dr. N.W., Calgary, Alberta T2N 1N4, Canada burke@math.ucalgary.ca D.M. Chibisov Steklov Mathematical Institute, Gubkina str. 8, 117966 Moscow, GSP-1, Russia chibisov@dmch.mian.su G. Christoph Otto-von-Guericke Universität, Fakultät für Mathematik, PF 4120, D-39016 Magdeburg, Germany Gerd.Christoph@Mathematik.Uni-Magdeburg.DE B. Courbot IRMAR, Université de Rennes 1, 35042 Rennes Cedex, France bernard.courbot@humana.univ-nantes.fr Csáki Endre MTA Rényi A. MKI, Budapest, Reáltanoda u. 13-15., P.O. Box 127, H - 1364, Hungary csaki@math-inst.hu Csiszár Imre MTA Rényi A. MKI, Budapest, Reáltanoda u. 13-15., P.O. Box 127, H - 1364, Hungary csiszar@math-inst.hu M. Csörgő School of Maths. Stats., Carleton University, 1125 Colonel By Drive, Ottawa, ON K1S 5B6, Canada mcsorgo@math.carleton.ca Csörgő Sándor JATE Bolyai Intézet, Szeged, Aradi vértanúk tere 1. H - 6720, Hungary csorgo@math.u-szeged.hu D.M. Dabrowska Department of Biostatistics, University of California, Los Angeles, CA 90095-1772, USA dorota@mozart.ph.ucla.edu Y. Davydov Lab. de Stat. et Probab., UFR de Mathématiques, Univ. Lille 1, 59655 Villeneuve d’Ascq Cedex, France davydov@jacta.univ-lille1.fr P. Deheuvels Dept. of Mathematics, Université Paris VI, 7 avenue du Chateau, 92340 Bourg-la-reine, France pd@ccr.jussieu.fr L. Devroye School of Comp. Sci., McGill University, 805 Sherbrooke ouest, Montreal, QC H3A 2K6, Canada luc@cs.McGill.CA S. Döhler Inst. f. Math. Stochastik, Albert-Ludwigs Univ. Freiburg, Eckerstrasse 1, D-79104 Freiburg, Germany doehler@stochastik.uni-freiburg.de G. Draisma Econometric Institute, Erasmus University Rotterdam, Room H11-12, P.O. Box 1838, 3000 DR Rotterdam, The Netherlands draisma@few.eur.nl E. Dyakonova Department of Discrete Mathematics, Steklov Mathematical Institute, Gubkin str. 8, 117966 Moscow GSP-1, Russia elena@mi.ras.ru N. Eisenbaum Lab. de Probabilités, Université Paris VI, 4 place Jussieu, 75252 Paris Cedex 05, France nae@ccr.jussieu.fr Ch. El-Nouty L.S.T.A. Univ. Paris VI, 4 Place Jussieu, 75252 Paris Cedex 05, France elnouty@ccr.jussieu.fr L. Ericksen P.O. Box 172, Millville, NJ 08332, USA le22@cornell.edu I. Fahrner Abteilung Mathematik III, Universität Ulm, D-89069 Ulm, Germany fahrner@mathematik.uni-ulm.de Fazekas István KLTE Matem. és Inform. Intézet, Debrecen, Pf. 12., H - 4010, Hungary fazekasi@math.klte.hu A. Földes Coll. of Staten Island, CUNY, Dept. of Statistics, 2800 Victory Blvd., Staten Island, NY 10314, USA afoldes@email.gc.cuny.edu K. Fukuyama Dept. of Math., Fac. Sci., Kobe University, Rokko, Kobe 675-8501, Japan fukuyama@math.kobe-u.ac.jp Gerencsér László MTA SZTAKI, Budapest, Kende u. 13-17., H - 1111, Hungary gerencser@sztaki.hu B. Gittenberger Department of Geometry, TU Wien, Wiedner Hauptstrasse 8-10/113, A-1040 Wien, Austria Bernhard.Gittenberger@tuwien.ac.at A. Gnedin University of Göttingen, Inst. of Math. Stochastics, Lotzestrasse 13, D-37083 Göttingen, Germany gnedin@math.uni-goettingen.de E. Gombay Univ. of Alberta, Dept. of Math. Sci., Edmonton, Alberta T6G 2G1, Canada gombay@stat.ualberta.ca Győrfi László BME Mat. Intézet, Budapest, Műegyetem rkp. 3-9., H - 1111, Hungary gyorfi@inf.bme.hu L. de Haan Erasmus University, Econometric Institute, P.O. Box 1738, 3000 DR Rotterdam, The Netherlands ldhaan@few.eur.nl K. Grill Inst. f. Statistik und, Wahrscheinlichkeitstheorie, TU Wien, Wiedner Hauptstrasse 8-10, A-1040 Wien, Austria grill@ci.tuwien.ac.at G. Haiman Université de Lille 1, 188 rue d’Alesia, 75014 Paris, France haiman@ccr.jussieu.fr D. Hlubinka KPMS MFF UK, Sokolovská 83, 186 75 Praha 8, Czech Republic hlubinka@karlin.mff.cuni.cz L. Horváth Department of Mathematics, University of Utah, 1440 East 155 South, Salt Lake City, UT 84112-0090, USA horvath@math.utah.edu Horváth Márta BME Mat. és Számítástud. Tsz., Budapest, Stoczek u. 2-4., H - 1111, Hungary marti@szit.bme.hu Yueyun Hu Laboratoire de Probabilités, Univ. Pierre et Marie Curie (PARIS 6), Tour 56-46, 4, place Jussieu, 75252 Paris Cedex 05, France hu@ccr.jussieu.fr M. Husková Charles University, Department of Statistics, Sokolovská 83, 186 75 Praha, Czech Republic huskova@karlin.mff.cuni.cz T. Inglot Inst. of Mathematics, Wroclaw University of Techn., Wybrzeze Wyspianskiego 27, 50-370 Wroclaw, Poland inglot@neyman.im.pwr.wroc.pl Ispány Márton KLTE Matem. és Inform. Intézet, Debrecen, Pf. 12., H - 4010, Hungary ispany@math.klte.hu M. Janžura Inst. of Inf. Theory and Autom., Czech Acad. of Sciences, P.O. Box 18, 182 08 Praha 8, Czech Republic janzura@utia.cas.cz A. Járai Cornell University, Department of Mathematics, Malott Hall, Ithaca, NY 14853, USA jarai@math.cornell.edu A. Jurlewicz Institute of Mathematics, Wroclaw Univ. of Technology, Wyb. wyspianskiego 27, 50-370 Wroclaw, Poland agniesz@im.pwr.wroc.pl H. Kaspi Cornell University, ORIE. 224 Rhodes Hall, Ithaca, NY 14853, USA kaspi@orie.cornell.edu V. Kolchin Steklov Mathematical Institute, Gubkin Str., 8, 117966 Moscow, GSP-1, Russia kolchin@mi.ras.ru A. Krzyżak Dept. of Computer Science, Concordia University, Montreal H3G 1M8, Canada KRZYZAK@CS.CONCORDIA.CA A. Lachal INSA DE LYON, Lab. de Modelis. Math. et Calc. Sci., Batiment 401, 24 Ave. Albert Einstein, 69621 Villeurbanne Cedex, France lachal@mathinsa.insa-lyon.fr P. Lachout KPMS MFF UK, Sokolovská 83, 186 75 Paha 8, Czech Republic lachout@karlin.mff.cuni.cz M.R. Leadbetter Dept. of Stat., Univ. of N.C., Chapel Hill, NC 27599, USA mrl@email.unc.edu J. Lember University of Tartu, Institute of Mathematical Statistics, Liivi 2-405, 50409 Tartu, Estonia jyril@madli.ut.ee E. Liebscher Technical University Ilmenau, Institute of Mathematics, PF 100565, D-98684 Ilmenau, Germany lieb@mathematik.tu-ilmenau.de H. Liero Institute of Mathematics, University of Potsdam, Am Neuen Palais 10, D-14415 Potsdam, Germany liero@rz.uni-potsdam.de M. Lifshits Komendantskii 22-2-49, 197372 St-Petersburg, Russia lifts@mail.rcom.ru T. Lin Erasmus Univ. of Rotterdam, H16-08, Fac. of Economics, P.O. Box 1738, 3000 DR Rotterdam, The Netherlands lin@few.eur.nl F. Lombard Dept. of Stats., Rand Afrikaans University, P.O. Box 524, 2006 Auckland Park, South Africa fred@na.rau.ac.za Major Péter MTA Rényi A. MKI, Budapest, Reáltanoda u. 13-15., P.O. Box 127, H - 1364, Hungary major@math-inst.hu Maricza István BME Távkl. és Telematikai Tsz.. Budapest, Pázmány P. sétány 1/D., H - 1117, Hungary maricza@cs.elte.hu Megyesi Zoltán JATE Bolyai Intézet, Szeged, Aradi vértanúk tere 1., H - 6720, Hungary megyesiz@server.math.u-szeged.hu Michaletzky György ELTE TTK Valószínűségelm. és Stat. Tsz., Budapest, Rákóczi út 5., H - 1088, Hungary michgy@ludens.elte.hu J. Mielniczuk Institute of Comp. Sci., Polish Academy of Sciences, Ordona 21, 01-237 Warsaw, Poland miel@ipipan.waw.pl Móri F. Tamás ELTE TTK Valószínűségelm. és Stat. Tsz., Budapest, Rákóczi út 5., H - 1088, Hungary moritamas@ludens.elte.hu P. Mörters Technische Universität Berlin, Fachbereich Mathematik Ma7-5, Strasse des 17. Juni 136, D-10623 Berlin, Germany peter@mathematik.uni-kl.de Morvai Gusztáv BME Mat. és Számítástud. Tsz., Budapest, Stoczek u. 2-4., H - 1111, Hungary morvai@inf.bme.hu N. Mushkudiani Dept. of Math. and Comp. Sci., Eindhoven Univ. of Technology, P.O. Box 513, 5600 MB Eindhoven, The Netherlands anino@win.tue.nl A. Nadas New York University, School of Medicine, 57 Old Forge Road, Tuxedo, NY 10987, USA nadas@charlotte.med.nyu.edu A.V. Nagaev Faculty of Mathematics, N. Copernicus University, 12/18 Chopin Str., 87-100 Torun, Poland nagaev@mat.uni.torun.pl G.L. O’Brien Dept. of Maths. and Stats., York University, 4700 Keele St., Toronto, Ontario M3J 1P3, Canada obrien@yorku.ca M. Orasch Technische Universität Wien, Institut für Statistik & Wahrscheinlichkeitstheorie, Wiedner Hauptstr. 8-10, A-1040 Wien, Austria orasch@ci.tuwien.ac.at S. Ostrovska Dokuz Eylul University, Faculty of Arts and Sciences, Department of Mathematics, Isciler Caddesi 143, 35230 Alsancak, Izmir, Turkey sofia@gediz.sci.deu.edu.tr Pap Gyula KLTE Mat. és Inf. Intézet, Debrecen, Pf. 12, H - 4010, Hungary papgy@math.klte.hu V. de la Peña Department of Statistics, Columbia University, New York, NY 10027, USA vp@stat.columbia.edu Petz Dénes BME Mat. Intézet, Budapest, Műegyetem rkp. 3-9., H - 1111, Hungary petz@math.bme.hu W. Philipp Dept. of Statistics, University of Illinois, 725 S. Wright Str., Champaign, IL 61820, USA philipp@stat.uiuc.edu A. Plucinska Mathematical Institute, Warsaw Technical University, pl. Politechniki 1, 00-644 Warsaw, Poland amp@alpha.im.pw.edu.pl G. Pritchard Dept. of Statistics, University of Auckland, Private Bag 92019, Auckland, New Zealand pritchar@scitec.auckland.ac.nz L. Rejtő Dept. of Math. Sci., Univ. of Delaware, Newark, DE 19711, USA rejto@udel.edu Révész Pál MTA Rényi A. MKI, Budapest, Reáltanoda u. 13-15., P.O. Box 127, H - 1364, Hungary reveszp@math-inst.hu L. Rüschendorf University of Freiburg, Inst. für Mathem. Stochastik, Eckerstr. 1, D-79104 Freiburg, Germany ruschen@stochastik.uni-freiburg.de B. Schmuland Dept. of Mathematical Sciences, University of Alberta, Edmonton, Alberta T6G 2G1, Canada schmu@stat.ualberta.ca J. Segers Departement Wiskunde, Celestijnenlaan 200B, B-3001 Heverlee, Belgium johan.segers@wis.kuleuven.ac.be Zhan Shi Université Paris VI, Laboratoire de Probabilités, 4 place Jussieu, 75252 Paris Cedex 05, France zhan@proba.jussieu.fr Gordon Simons Department of Statistics, Univ. of North Carolina, Chapel Hill, NC 27599-3260, USA simons@stat.unc.edu A. Spătaru Casa Academiei Romane, Calea 13 Septembrie nr. 13, 76100 Bucharest 5, Romania aspataru@pcnet.pcnet.ro J.G. Steinebach Philipps Universität, FB Mathematik and Informatik, Hans-Meerwein-Str., D-35032 Marburg, Germany jost@mathematik.uni-marburg.de Ch. Suquet Lab. de Stat. et Probab., Bât. M2, Cité Scientifique, 59655 Villeneuve d’Ascq Cedex, France Charles.Suquet@univ-lille1.fr Szabados Tamás BME Mat. Intézet, Budapest, Műegyetem rkp. 3-9., H - 1111, Hungary szabados@math.bme.hu P. Szabłowski Institute of Mathematics, Technical Univ. of Warsaw, Pl. Politechniki 1, 00-661 Warsaw, Poland p.szablowski@elka.pw.edu.pl Szász Domokos MTA Rényi A. MKI, Budapest, Reáltanoda u. 13-15., P.O. Box 127, H - 1364, Hungary szasz@math-inst.hu G.J. Székely Dept. of Math., Bowling Green State Univ., Bowling Green, OH 43403-0221, USA gabors@math.bgsu.edu Z.S. Szewczak Nicholas Copernicus Univ., NTC, ul. Chopina 12/18, 87-100 Torun, Poland zssz@mat.uni.torun.pl S.J. Taylor 6 Winters Lane, Portesham, Dorset DT3 4ES, United Kingdom sjt4k@aol.com Terdik György KLTE Inf. és Számítókp., Debrecen, Pf. 58., H - 4010, Hungary Terdik@CIC.klte.hu E. Thilly Lab. de Stat. et Probab., Université de Lille 1, 59655 Villeneuve d’Ascq, France thilly@jacta.univ-lille1.fr E. Ursianu Casa Academiei Romane, Calea 13 Septemberie nr. 13, 76100 Bucharest 5, Romania eursianu@ns.csm.ro I. Van Keilegom Penn State University, Department of Statistics, 413 Thomas Building, University Park, PA 16802, USA ivankeil@stat.psu.edu K. Van Zyl Dept. of Statistics, Rand Afrikaans University, P.O. Box 524, 2006 Auckland Park, South Africa jjvz@na.rau.ac.za Viharos László JATE Bolyai Intézet, Szeged, Aradi vértanúk tere 1., H - 6720, Hungary viharos@math.u-szeged.hu D. de Waal Dept. Math. Statistics, Univ. of the Orange Free State, P.O. Box 339, Bloemfontein 9300, South Africa wwdw@wwg3.uovs.ac.za K. Yamada Department of Mathematics, Kanagawa University, Kanagawa-ku, Yokohama, 221, Japan keigo@cc.kanagawa-u.ac.jp M. Yor Lab. de Probabilités, UMR 7599, Université Paris VI, 4 place Jussieu, Tour 56, 75252 Paris Cedex 05, France secret@proba.jussieu.fr V. Zaiats Dept. Matemàtica i Fisíca Apl., Universitat de Vic, c/. Sagrada Família, 7, 08500 Vic (Barcelona), Spain ZAIATS@MAT.UAB.ES O. Zaihraiev N. Copernicus University, Fac. of Math. and Informatics, Chopin str. 12/18, 87-100 Torun, Poland alzaig@mat.uni.torun.pl A. Zaitsev St. Petersburg Branch of Steklov Math. Inst., Fontanka 27, 191011 St. Petersburg, Russia zaitsev@pdmi.ras.ru R. Zitikis Department of Statistics, University of Manitoba, Winnipeg, Manitoba R3T 2N2, Canada Ricardas Zitikis@umanitoba.ca